OpenStudy (anonymous):

find the equation of the line having the given slope and containing the given point m=6/7,(8,-3) PLEASEEEE HELP

OpenStudy (anonymous):

y-y0 = m(x - x0) this is the formula you use. you put 6/7 for m, and you put x0: 8 y0: -3 then, expand the parentheses and bring into standard form

OpenStudy (anonymous):

this is what i got y+3=6/7(x-8)

OpenStudy (anonymous):

very good try to expand the parentheses

OpenStudy (anonymous):

hmm y+3=6/7x and then do i multiply 6/7 by -8 ? its kind confusing

OpenStudy (anonymous):

it is difficult. and yes, you multiply 6/7 with -8 when multiplying you just multiply the upper term, so 6 x -8 = -48 for the sign: negative times positive = negative we just carry over the denominator when we are done and get: -48 / 7

OpenStudy (anonymous):

so is it y+3=6/7x-48/7 ? or do i subtract 3 from both sides ?

OpenStudy (anonymous):

you also subtract three from both sides as the final step :) then it's in standard form y=mx +b

OpenStudy (anonymous):

so its y=6/7x-51/7 is the final answer? and thank you so much (:

OpenStudy (anonymous):

something like that .if you are subtracting a number from a fraction you must first make that number have the same denominator

OpenStudy (anonymous):

i dont get it ;(

OpenStudy (anonymous):

applying the operation to the upper term only works for multiplication exlusively in case of add/subtract: \[\frac{-48}{7} - 3 =\frac{-48}{7} - \frac{3 \times 7}{7}\]\[\frac{-48}{7} - \frac{21}{7} = \frac{-69}{7}\]

OpenStudy (anonymous):

in other words you need the same denominator for addition/subtraction!!

OpenStudy (anonymous):

so i put -69/7 instead instead of -51/7 ? thats it?

OpenStudy (anonymous):

yes you can extend "3" into "21/7" and then subtract the 21/7 from what we had an you would get the same result

OpenStudy (anonymous):

3/4 - 1/4 = 2/4 <- this is a rule 3/4 - 1 = <- this needs the first denominator first 3/4 -4/4 = -1/4 <- aha!

OpenStudy (anonymous):

you cant do everything at the same time this is our result y=6/7x-69/7

OpenStudy (anonymous):

so what is the final answer ? cause im lost D:

OpenStudy (anonymous):

y=6/7x-69/7

OpenStudy (anonymous):

okay thank you alotttt (: c: :3

OpenStudy (anonymous):

np